Can you please help me sove this?

2x-3y=1.3, y-x=0.5
1
Expert's answer
2014-07-21T13:54:21-0400
From the second equation of the system we express y=x+0.5 and plug into the first equation and obtain
2x-3y=1.3, 2x-3(x+0.5)=1.3, simplify -x-1.5=1.3, whence x=-1.5-1.3, therefore, x=-2.8. Recall y=x+0.5, therefore y=-2.8+0.5, y=-2.3.
Answer: x=-2.8, y=-2.3.